### v2

# Probing $\alpha$ clustering in $^{12}\mathrm{C}$ at CSR energies using the Jet AA Microscopic Transport Model

Abstract
We investigate the sensitivity of low-energy nuclear collisions to intrinsic nuclear structure by studying the interplay between initial-state geometry and final-state observables in C+C and C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV, relevant for experiments at the Cooling Storage Ring (CSR) facility in Lanzhou and forthcoming experiments at the High Intensity heavy-ion Accelerator Facility (HIAF) in Huizhou. Calculations are performed within the Jet AA Microscopic Transport Model (JAM) using Woods–Saxon and triangular $\alpha$-clustered configurations for the $^{12}\mathrm{C}$ nucleus. The initial geometry is characterized in terms of transverse size, compactness, eccentricities, and their ensemble-averaged fluctuations. We find that $\alpha$ clustering leads to a more compact participant configuration than the Woods–Saxon case, while transverse-size and eccentricity fluctuations show only weak sensitivity to clustering. At this beam energy, radial observables remain sensitive to geometric compactness, with the ensemble-averaged proton mean transverse momentum $\langle p_{T}\rangle$ enhanced for $\alpha$-clustered configurations, whereas pions show little sensitivity. The anisotropic response is examined using flow harmonic coefficients. We find an enhancement of the root-mean-square flow magnitudes, $v_{n}\{2\}=\sqrt{\langle v_{n}^{2}\rangle}$, for $\alpha$-clustered configurations at large $N_{\mathrm{part}}$, while the ensemble-averaged fluctuation strength of individual harmonics remains small. Symmetric cumulants of the initial-state eccentricities show sensitivity to clustering, whereas the corresponding ensemble-averaged correlations among final-state flow harmonics do not exhibit a comparably strong separation. These results indicate that radial observables and correlation-based flow measurements provide complementary probes of $\alpha$ clustering in low-energy nuclear collisions.
I Introduction
In recent years, a new direction in relativistic heavy-ion physics has emerged that exploits the geometric and structural properties of the colliding nuclei as tools to probe both nuclear structure and the collective response of the produced medium. Experimental measurements have demonstrated that intrinsic nuclear deformation can leave observable imprints on final-state momentum anisotropies. In particular, the STAR Collaboration has reported clear signatures of the quadrupole deformation of uranium nuclei in U+U collisions at RHIC energies M. I. Abdulhamid et al. (2024); 23, while measurements in Xe+Xe collisions at the LHC Acharya and others (2018); Sirunyan and others (2019); Aad and others (2020) have revealed sensitivity to the prolate deformation of the xenon nucleus through deviations from expectations based on spherical nuclei.
These developments have motivated growing interest in using light nuclei with pronounced intrinsic structure as controlled probes of initial-state geometry in nuclear collisions. In this context, collisions involving nuclei such as $^{12}\mathrm{C}$ and $^{16}\mathrm{O}$ are particularly intriguing, as they are predicted to exhibit strong $\alpha$-cluster correlations in their ground-state wave functions Broniowski and Ruiz Arriola (2014); He et al. (2014); Rybczyński et al. (2018); Zhang et al. (2017); Guo et al. (2017); Yao et al. (2025); Ma et al. (2020). The intrinsic structure of $^{12}\mathrm{C}$, and in particular the nature of its Hoyle state, has been the subject of extensive theoretical and experimental investigations. While ab initio calculations based on effective field theory have suggested that the Hoyle state may resemble a linear chain of three $\alpha$ clusters Epelbaum et al. (2011), more recent experimental measurements favor a predominantly triangular $\alpha$-cluster configuration Marin-Lambarri et al. (2014); Shen et al. (2023), indicating that its underlying geometry remains an active subject of study.
Recent experimental results from O+O collisions at RHIC 11 and the LHC Collaboration and others (2025); Acharya and others (2025); Aad and others (2025, 2026); Hayrapetyan and others (2025), together with complementary theoretical studies, suggest that small-system collisions may retain sensitivity to such subnucleonic or cluster-scale geometric features. These observations open the possibility of accessing clustering effects through final-state observables in relativistic nuclear collisions. Since $\alpha$ clustering is an intrinsic feature of the initial coordinate-space geometry of the nucleus, its experimental manifestation necessarily relies on the efficiency with which these spatial correlations are dynamically translated into measurable momentum-space signatures.
At ultra-relativistic collision energies, the connection between nuclear geometry and final-state observables is commonly interpreted within a hydrodynamic framework. In this picture, initial-state spatial anisotropies and density inhomogeneities are efficiently converted into momentum-space correlations through the collective expansion of a nearly equilibrated medium. The system’s response is predominantly governed by pressure gradients and transport properties, enabling nuclear structure effects to be investigated in a controlled manner via their influence on initial eccentricities and density profiles.
At lower collision energies, however, the dynamical evolution of the system is qualitatively different. Baryon stopping, finite net-baryon density, mean-field potentials, and hadronic rescattering play a central role and can substantially modify the mapping between the initial geometry and final-state observables. Consequently, it is not a priori evident whether, or to what extent, nuclear structure effects such as $\alpha$ clustering survive the subsequent evolution and remain observable. Determining how initial geometric features are preserved in this regime therefore constitutes an important open question and is essential for assessing the sensitivity of low-energy heavy-ion collisions to nuclear structure.
Recent and forthcoming experimental programs provide new opportunities to explore these questions. Experiments involving C+C and C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV have recently commenced at the Cooling Storage Ring (CSR) facility in Lanzhou  et al. (2017); Xia and others (2002); Mao and others (2020); Zhu et al. (2021); Liu et al. (2023); Sahoo et al. (2024); Hu et al. (2023); Wu et al. (2026). In addition, the upcoming High Intensity heavy-ion Accelerator Facility (HIAF) J. C. Yang et al. (2013); X. Zhou and J. Yang (2022); J. Yang (2025); 19 in Huizhou will enable similar studies with a variety of collision systems, including O+O and U+U. Significant efforts are also underway at the LHC, where collisions of nuclei with non-spherical or exotic configurations, such as deformed Ne+Ne systems Giacalone and others (2025); Aad and others (2025), have been performed recently. Together, these programs open a new avenue for systematically investigating nuclear structure effects across a wide range of collision energies.
Figure 1: Radial distribution of the point-proton density of $^{12}\mathrm{C}$ obtained from JAM initial configurations for Woods–Saxon and $\alpha$-clustered nuclei. The densities are normalized to the total proton number.
II JAM Framework
In this work, we employ the Jet AA Microscopic Transport Model (JAM), a hadronic transport framework developed for the simulation of relativistic nuclear collisions over a broad range of beam energies Nara et al. (2000); Hirano:2012yy; Hirano et al. (2013); Isse et al. (2005); Nara and Stoecker (2019); Nara et al. (2020); Nara and Ohnishi (2022). In JAM, the initial positions of nucleons inside each nucleus are sampled according to a prescribed nuclear density distribution, after which the collision dynamics are modeled as a sequence of binary hadron–hadron interactions combined with mean-field effects. The model describes the full time evolution of the system, from the initial nuclear overlap through the dense interaction stage to the final hadronic freeze-out.
For the present study, JAM (version 2.58) Nara and others is used in its mean-field mode, which incorporates nuclear equation-of-state effects through a momentum-dependent potential that modifies the propagation of hadrons. This approach is essential at intermediate collision energies, where baryon stopping and mean-field interactions play a significant role in shaping collective dynamics. Events are generated for C+C and C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV, both with and without $\alpha$ clustering in the $^{12}\mathrm{C}$ nucleus, where the implementation of clustering will be discussed in detail in the following section. We use the MD2 equation of state Nara et al. (2020), characterized by a nuclear incompressibility of $K=380$ MeV, which has been shown to successfully reproduce collective flow observables in low-energy heavy-ion collisions at RHIC and lower energies. The time evolution is followed up to $100~\mathrm{fm}/c$ with a time step of $1~\mathrm{fm}/c$.
II.1 Clustering implementation
To investigate the role of nuclear structure effects, two distinct initial configurations are employed for the $^{12}\mathrm{C}$ nucleus. In the reference configuration, nucleon positions are sampled from a spherical Woods–Saxon distribution. In the alternative configuration, $^{12}\mathrm{C}$ is modeled as an $\alpha$-clustered system consisting of three $\alpha$ clusters arranged at the vertices of an equilateral triangle in the transverse plane. This triangular configuration is motivated by nuclear structure studies, which have long predicted a pronounced three-$\alpha$ cluster structure in the ground state of $^{12}\mathrm{C}$ Hafstad and Teller (1938); Harrington (1966); Ikeda et al. (1968); Noble (1970); Bertsch and Bertozzi (1971). Both transport and hydrodynamic models have been extensively employed to explore possible experimental signatures of nuclear clustering in heavy-ion collisions over a wide range of beam energies Broniowski and Ruiz Arriola (2014); He et al. (2014); Rybczyński et al. (2018); Zhang et al. (2017); Guo et al. (2017); Ma et al. (2020); Liu et al. (2024); Yao et al. (2025).
The centers of the $\alpha$ clusters are separated by a fixed distance $d_{\alpha\alpha}$, which sets the characteristic length scale of the clustered geometry. Within each $\alpha$ cluster, the four constituent nucleons are distributed around the corresponding cluster center according to a Gaussian profile, $\rho_{\alpha}(\mathbf{r})\propto\exp\!\left(-\mathbf{r}^{2}/2\sigma^{2}\right)$, where the width parameter $\sigma$ controls the spatial localization of nucleons inside the cluster. In the present implementation, the values of $d_{\alpha\alpha}$ and $\sigma$ are fixed for all events, while the overall orientation of the triangular configuration is randomized on an event-by-event basis. To prevent unphysical overlap between nucleons, a hard-core distance is imposed between all nucleon pairs.
The parameters of the $\alpha$-clustered configuration ($d_{\alpha\alpha}$=3.1 fm and $\sigma$=0.8) are chosen such that the resulting point-proton density distribution and root-mean-square radius of $^{12}\mathrm{C}$ remain comparable to those obtained from the Woods–Saxon configuration. This ensures that the two initial conditions differ primarily in their internal spatial organization and short-range correlations, rather than in their global size or average density, thereby enabling a controlled study of clustering-induced geometric effects. With this setup, JAM provides a consistent framework in which the impact of $\alpha$ clustering on the initial geometry and its subsequent influence on final-state observables can be systematically investigated while keeping the underlying transport dynamics and equation of state fixed.
The proton density distribution of $^{12}\mathrm{C}$ is constructed from the initial-state nucleon configurations generated in JAM by accumulating proton coordinates over all events to obtain a spherically averaged density,
$\rho_{p}(r)=\frac{1}{4\pi r^{2}}\frac{dN_{p}}{dr},$ (1)
which is normalized according to
$\int d^{3}r\,\rho_{p}(r)=6.$ (2)
No additional smearing associated with the finite proton charge radius is applied, such that the resulting distribution corresponds to the point-proton density. The same procedure is used consistently for both Woods–Saxon and $\alpha$-clustered initial configurations.
Differences between the two configurations arise primarily from their internal spatial structure rather than from the global density scale. In particular, $\alpha$ clustering modifies short-range correlations and local proton packing while leaving the angle-averaged density largely unchanged. This is reflected in the fact that the root-mean-square (RMS) radius of the Woods–Saxon configuration of $^{12}\mathrm{C}$ is $r_{\mathrm{rms}}\simeq 2.4$ fm, while that of the $\alpha$-clustered configuration is $r_{\mathrm{rms}}\simeq 2.3$ fm, corresponding to a difference of less than $4\%$. This demonstrates that the observed differences in initial-state compactness and in the final-state proton mean transverse momentum do not originate from trivial changes in the overall nuclear size or charge density, but instead reflect differences in the microscopic spatial organization of the nucleons.
Figure 2: The transverse size ($\langle r^{2}\rangle$), compactness ($C$), and the corresponding transverse-size fluctuation strength as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+C collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
Figure 3: The transverse size ($\langle r^{2}\rangle$), compactness ($C$), and the corresponding transverse-size fluctuation strength as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
III Results and discussions
III.1 Correlation in initial-state geometries
To characterize the geometric properties of the initial participant configuration, we consider several complementary observables constructed from the transverse coordinates of participant nucleons. The average transverse size is quantified by the second moment
$\langle r^{2}\rangle=\frac{1}{N_{\mathrm{part}}}\sum_{i=1}^{N_{\mathrm{part}}}\left[(x_{i}-\bar{x})^{2}+(y_{i}-\bar{y})^{2}\right],$ (3)
where $(\bar{x},\bar{y})$ denotes the transverse center of mass of the participants in a given event. To remove the trivial scaling with the number of participants, we also study the scaled transverse size $\langle r^{2}\rangle/N_{\mathrm{part}}$. The corresponding compactness of the participant zone is defined as
$C\equiv\sqrt{\frac{N_{\mathrm{part}}}{\langle r^{2}\rangle}},$ (4)
which provides a measure of the inverse transverse length scale governing the initial pressure gradients. Fluctuations of the transverse size are quantified by the relative width
$\frac{\sigma(\langle r^{2}\rangle)}{\langle r^{2}\rangle},$ (5)
evaluated as an ensemble average at fixed $N_{\mathrm{part}}$.
Figure 2 and 3 present the scaled transverse size ($\langle r^{2}\rangle/N_{\mathrm{part}}$), compactness ($C$), and the corresponding transverse-size fluctuation strength as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+C and C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV. In C+C collisions, the $\alpha$-clustered configuration exhibits a systematically smaller scaled transverse size than the Woods–Saxon case, indicating a more compact spatial arrangement of the participant nucleons. This reduction in $\langle r^{2}\rangle/N_{\mathrm{part}}$ is accompanied by a corresponding enhancement of the compactness over the full $N_{\mathrm{part}}$ range. In contrast, the relative transverse-size fluctuations show a weaker sensitivity to clustering, with comparable magnitudes for Woods–Saxon and $\alpha$-clustered configurations, while the magnitude of the fluctuations decreases with increasing $N_{\mathrm{part}}$. The observed pattern demonstrates that, in symmetric light–light collisions, $\alpha$ clustering primarily shifts the average transverse scale of the system rather than significantly enhancing transverse-size fluctuations.
A different behavior is observed in C+Pb collisions. While the average compactness remains larger for the $\alpha$-clustered configuration, reflecting the increased local packing of nucleons in the carbon projectile, the relative transverse-size fluctuations are comparable or slightly larger for the Woods–Saxon case. In this asymmetric system, fluctuations of the transverse size are dominated by the geometry and participant fluctuations on the Pb side, which reduces the relative impact of clustering in the lighter nucleus. Consequently, $\alpha$ clustering continues to modify the mean transverse geometry but does not lead to an enhancement of transverse-size fluctuations. Overall, the comparison shown in Figs. 2 and 3 demonstrates that $\alpha$ clustering predominantly affects the average compactness of the participant zone, while its influence on transverse-size fluctuations is comparatively modest and most pronounced at low $N_{\mathrm{part}}$.
Figure 4: Ensemble-averaged participant eccentricities and their corresponding fluctuation strengths as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+C collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
In addition to transverse size observables, the anisotropic geometry of the initial participant configuration is usually characterized using event-by-event eccentricities $\varepsilon_{n}$. For each event, participant nucleons are identified and their transverse coordinates $(x_{i},y_{i})$ are evaluated relative to the participant center of mass. The eccentricity coefficients are defined using weighted moments of the participant distribution Alver and Roland (2010),
$\varepsilon_{n}=\frac{\sqrt{\left\langle w_{n}\cos(n\phi)\right\rangle^{2}+\left\langle w_{n}\sin(n\phi)\right\rangle^{2}}}{\left\langle w_{n}\right\rangle},$ (6)
where $\phi=\tan^{-1}(y/x)$ is the azimuthal angle of a participant. For the dipole eccentricity ($n=1$), the weight $w_{1}=r^{3}$ is used, while for the elliptic ($n=2$) and triangular ($n=3$) eccentricities the weight $w_{2,3}=r^{2}$ is employed. These definitions follow standard prescriptions in studies of initial-state geometry and ensure sensitivity to long-wavelength shape deformations.
Event-averaged eccentricities are obtained as functions of $N_{\mathrm{part}}$ by averaging $\varepsilon_{n}$ over events within a given $N_{\mathrm{part}}$ bin. The fluctuation strength is quantified using the variance
$\sigma(\varepsilon_{n})=\sqrt{\langle\varepsilon_{n}^{2}\rangle-\langle\varepsilon_{n}\rangle^{2}},$ (7)
as well as the corresponding relative fluctuation $\sigma(\varepsilon_{n})/\langle\varepsilon_{n}\rangle$. These quantities characterize the intrinsic geometric variability of the initial state at fixed system size.
Figure 4 presents the average eccentricities $\varepsilon_{n}$ and their corresponding fluctuation strengths as a function of $N_{\mathrm{part}}$ for C+C collisions. For both the Woods–Saxon and $\alpha$-clustered configurations, the eccentricities follow the ordering $\varepsilon_{2}>\varepsilon_{3}>\varepsilon_{1}$, with comparable magnitudes in the two cases. The relative fluctuations of each $\varepsilon_{n}$ remain approximately constant as a function of $N_{\mathrm{part}}$, with the ordering $\sigma(\varepsilon_{1})>\sigma(\varepsilon_{3})>\sigma(\varepsilon_{2})$.
Figure 5: Ensemble-averaged participant eccentricities and their corresponding fluctuation strengths as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV. Results are calculated using C participants only (see text for details)
Figure 6: Ensemble-averaged participant eccentricities and their corresponding fluctuation strengths as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
For C+Pb collisions, the eccentricities are calculated in two distinct ways. First, to isolate the effect of the carbon projectile, we compute $\varepsilon_{n}$ using only participants from the $^{12}\mathrm{C}$ nucleus; these results are shown in Fig. 5. In the $\alpha$-clustering case, $\varepsilon_{2}$ and $\varepsilon_{3}$ exhibit slighly opposite trends with $N_{\mathrm{part}}$: one increases while the other decreases, and a crossing near $N_{\mathrm{part}}\sim 60$. In contrast, for the Woods–Saxon configuration both $\varepsilon_{2}$ and $\varepsilon_{3}$ decrease with $N_{\mathrm{part}}$. A similar pattern is observed in the relative fluctuations of $\varepsilon_{n}$.
Second, we compute the eccentricities using participants from both the Pb and C nuclei, which is the more natural choice for the full collision system. These results are presented in Fig. 6. It is important to note that in both Figs. 5 and 6, the horizontal axis represents the total $N_{\mathrm{part}}$ from both nuclei. Although the C-only participant calculation showed clear differences between Woods–Saxon and $\alpha$-clustering configurations, these differences become less obvious when participants from both target and projectile are included.
Figure 7: Symmetric cumulants of eccentricties as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+C collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV
Figure 8: Symmetric cumulants of eccentricties as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV. Results are calculated using C participants only (see text for details)
Figure 9: Symmetric cumulants of eccentricties as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV. Results are calculated using C participants only (see text for details)
Correlations between different harmonic components of the initial geometry are further investigated using mixed moments such as $\langle\varepsilon_{1}\varepsilon_{3}\rangle$ and $\langle\varepsilon_{2}\varepsilon_{3}\rangle$, as well as their squared counterparts. These observables probe the degree to which different geometric deformations are correlated within an ensemble of events and provide information beyond that contained in individual eccentricity magnitudes.
To quantify such correlations in a more systematic manner, we evaluate symmetric cumulants constructed from the eccentricities. For two harmonics $m$ and $n$, the symmetric cumulant is defined as
$\mathrm{SC}\{m,n\}=\langle\varepsilon_{m}^{2}\,\varepsilon_{n}^{2}\rangle-\langle\varepsilon_{m}^{2}\rangle\langle\varepsilon_{n}^{2}\rangle,$ (8)
where the angular brackets denote ensemble averages over events at fixed $N_{\mathrm{part}}$. Positive values of $\mathrm{SC}\{m,n\}$ indicate a positive correlation between $\varepsilon_{m}^{2}$ and $\varepsilon_{n}^{2}$, while negative values signal an anticorrelation. In this work, symmetric cumulants are evaluated for the harmonic combinations $(m,n)=(1,3)$ and $(2,3)$ as functions of $N_{\mathrm{part}}$. For presentation purposes, the cumulants are scaled by $N_{\mathrm{part}}$ to reduce the trivial dilution of correlations with increasing system size.
The correlations among the eccentricities $\mathrm{SC}\{\varepsilon_{1},\varepsilon_{3}\}$ and $\mathrm{SC}\{\varepsilon_{2},\varepsilon_{3}\}$ are presented in Figs. 7, 8, and 9 for C+C collisions, C+Pb collisions with C participants only, and C+Pb collisions with all participants, respectively. In C+C collisions, no obvious differences are observed between the $\alpha$-clustered and Woods–Saxon configurations. For C+Pb collisions when considering C participants only, $\mathrm{SC}\{\varepsilon_{1},\varepsilon_{3}\}$ in the Woods–Saxon case exhibits a positive correlation that increases with $N_{\mathrm{part}}$, whereas the $\alpha$-clustered case shows an opposite trend. For $\mathrm{SC}\{\varepsilon_{2},\varepsilon_{3}\}$, the Woods–Saxon configuration remains nearly constant around zero as a function of $N_{\mathrm{part}}$, while the clustered configuration displays an increasingly negative correlation with increasing $N_{\mathrm{part}}$. Similarly, when all participants are included, $\mathrm{SC}\{\varepsilon_{1},\varepsilon_{3}\}$ shows a positive correlation for both Woods–Saxon and clustered configurations, although the magnitude in the clustered case is smaller than that in the Woods–Saxon case. For $\mathrm{SC}\{\varepsilon_{2},\varepsilon_{3}\}$, the qualitative features are similar to those observed in the C-participants-only case, but with reduced magnitudes. These results indicate that $\alpha$ clustering can induce a negative correlation between $\varepsilon_{2}$ and $\varepsilon_{3}$, which may be reflected in the final-state flow coefficients $v_{2}$ and $v_{3}$, whereas the correlation between $v_{1}$ and $v_{3}$ may not be easily distinguishable between the two configurations.
III.2 Bulk observables with final-state particles
Figure 10: $dN/d\eta$ of protons and pions for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+C collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
Figure 11: $dN/d\eta$ of protons and pions for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
All observables defined in previous section are constructed solely from the initial-state nucleon positions and thus directly reflect the geometric properties of the participant configuration. The impact of these initial-state features on the collision dynamics is subsequently assessed through final-state observables, including the yield, mean transverse momentum of identified particles, flow coefficients $v_{n}$, and correlations between $v_{n}$ and symmetric cumulants. This combined analysis allows for a systematic investigation of how initial-state geometry and clustering effects are transmitted to experimentally accessible final-state observables.
Figure 12: $\langle p_{T}\rangle$ versus $\eta$ of protons and pions for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+C collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
Figure 13: $\langle p_{T}\rangle$ versus $\eta$ of protons and pions for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
First, the bulk observables from produced particles are examined to assess the influence of $\alpha$-clustering on final-state observables and their response to the initial-state geometry. Pseudorapidity distributions $dN/d\eta$ for protons and pions are shown in Fig. 10 for C+C collisions and in Fig. 11 for C+Pb collisions, with results presented for both Woods–Saxon and $\alpha$-clustering configurations. The corresponding average transverse momentum $\langle p_{T}\rangle$ distributions are displayed in Fig. 12 and Fig. 13. In C+C collisions, $\langle p_{T}\rangle$ is symmetric about mid-rapidity, as expected for symmetric systems. In asymmetric C+Pb collisions, $\langle p_{T}\rangle$ distributions exhibit the expected asymmetry. When integrated over $N_{\mathrm{part}}$, the bulk observables are found to be similar for both initial-state configurations.
III.3 Correlation in final-state particles
The observed hierarchy in participant-zone compactness and transverse-size fluctuations provides a natural geometric explanation for the stronger sensitivity of mean radial observables, such as the proton $\langle p_{T}\rangle$, to clustering effects compared to fluctuation-driven observables.
The event-wise mean transverse momentum of identified particles is defined as
$\langle p_{T}\rangle_{\alpha}=\frac{1}{M_{\alpha}}\sum_{i=1}^{M_{\alpha}}p_{T,i},$ (9)
where $\alpha$ denotes the particle species and $M_{\alpha}$ is the corresponding multiplicity in a given event. Event-ensemble averages are obtained by further averaging $\langle p_{T}\rangle_{\alpha}$ over events at fixed $N_{\mathrm{part}}$, which we denote by double angle brackets.
Transverse-momentum fluctuations are quantified using the variance of the single-particle transverse-momentum distribution within a fixed $N_{\mathrm{part}}$ class,
$\delta p_{T,\alpha}^{2}=\langle p_{T}^{2}\rangle_{\alpha}-\langle p_{T}\rangle_{\alpha}^{2},$ (10)
where the averages are taken over all particles from all events belonging to the given $N_{\mathrm{part}}$ bin.
To probe relative changes between particle species, we also study the event-wise difference between the proton and pion mean transverse momenta,
$\Delta p_{T}^{(p-\pi)}=\langle p_{T}\rangle_{p}-\langle p_{T}\rangle_{\pi},$ (11)
Figures 14 and 15 show $\langle p_{T}\rangle$ for protons and pions as a function of $N_{\mathrm{part}}$ in C+C and C+Pb collisions, respectively. In both collision systems, protons exhibit a larger $\langle p_{T}\rangle$ in the $\alpha$-clustered configuration than in the Woods–Saxon case, whereas pions show no significant difference between the two configurations. This pattern is consistent with expectations based on the hierarchy of participant-zone compactness.
The difference in $\langle p_{T}\rangle$ between protons and pions is also shown. A clear distinction is observed in C+C collisions, while in C+Pb collisions the effect is most pronounced in the mid-$N_{\mathrm{part}}$ region.
Figure 14: $\langle p_{T}\rangle$ of protons and pions and their differences as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+C collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
Figure 15: $\langle p_{T}\rangle$ of protons and pions and their differences as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
To investigate the response of transverse-momentum fluctuations for different particle species, the variance of the single-particle transverse-momentum distribution is evaluated separately for each species $i$ within a fixed $N_{\mathrm{part}}$ class,
$\delta p_{T,i}^{2}=\Big\langle\!\Big\langle\langle p_{T}^{2}\rangle_{i}-\langle p_{T}\rangle_{i}^{2}\Big\rangle\!\Big\rangle.$ (12)
Here, $\langle\cdots\rangle_{i}$ denotes an average over all particles of species $i$ and over all events belonging to the given $N_{\mathrm{part}}$ bin. The index $i$ labels the particle species (e.g. $\pi$ or $p$).
The results for C+C and C+Pb collisions are shown in Figs. 16 and 17, respectively. In both collision systems, the $\langle p_{T}\rangle$ fluctuations of protons are smaller in the $\alpha$-clustered configuration than in the Woods–Saxon case, consistent with the reduced transverse-radius fluctuations of participant nucleons in the clustered configuration.
Figure 16: Ensemble-averaged single-particle transverse-momentum variance $\delta p_{T,i}^{2}$ as a function of $N_{\mathrm{part}}$ for different particle species in C+C collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
Figure 17: Ensemble-averaged single-particle transverse-momentum variance $\delta p_{T,i}^{2}$ as a function of $N_{\mathrm{part}}$ for different particle species in C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
To characterize the magnitude and event-by-event fluctuations of the flow harmonics, we evaluate moments of the event-wise flow coefficients using the $Q$-vector formalism. For each event, the quantity $v_{n}^{2}=|Q_{n}|^{2}/M^{2}$ is constructed, where $Q_{n}=\sum_{i}e^{in\phi_{i}}$ and $M$ denotes the multiplicity of selected particles. The typical magnitude of the flow harmonics is characterized by the root-mean-square value $v_{n}\{2\}=\sqrt{\langle v_{n}^{2}\rangle}$, while the fluctuation strength of a given harmonic is quantified by the variance $\langle v_{n}^{4}\rangle-\langle v_{n}^{2}\rangle^{2}$, with angle brackets denoting an average over events at fixed $N_{\mathrm{part}}$. The quantity $v_{n}\{2\}$ corresponds to the root-mean-square flow coefficient and is sensitive to both the mean flow and its event-by-event fluctuations. Only events with charged-particle multiplicity $N_{\mathrm{ch}}(|\eta|<2)>4$ are included in the analysis.
Figure 18: Root-mean-square flow magnitude $v_{1}\{2\}=\sqrt{\langle v_{1}^{2}\rangle}$ and the corresponding variance $\langle v_{1}^{4}\rangle-\langle v_{1}^{2}\rangle^{2}$ as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
Figure 19: Root-mean-square flow magnitude $v_{1}\{2\}=\sqrt{\langle v_{2}^{2}\rangle}$ and the corresponding variance $\langle v_{2}^{4}\rangle-\langle v_{2}^{2}\rangle^{2}$ as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
Figure 20: Root-mean-square flow magnitude $v_{3}\{2\}=\sqrt{\langle v_{3}^{2}\rangle}$ and the corresponding variance $\langle v_{3}^{4}\rangle-\langle v_{3}^{2}\rangle^{2}$ as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
Figures 18, 19, and 20 present the root-mean-square (RMS) flow magnitude and the corresponding fluctuation strength for the first-, second-, and third-order flow harmonics, evaluated using ensemble-averaged moments at fixed $N_{\mathrm{part}}$. We observe that $\alpha$ clustering leads to a significant enhancement of the RMS flow magnitude, particularly at large $N_{\mathrm{part}}$ ($\gtrsim 70$). In contrast, the variance $\langle v_{n}^{4}\rangle-\langle v_{n}^{2}\rangle^{2}$ of individual harmonics remains small and is not clearly distinguishable within the current statistical precision. This indicates that $\alpha$ clustering primarily modifies the average initial geometry and its collective response, while inducing only modest changes in the magnitude of single-harmonic flow fluctuations. Consequently, precise experimental measurements of flow observables in this regime could provide valuable constraints on the presence and strength of $\alpha$ clustering in light nuclei.
Finally, ensemble-averaged symmetric cumulants of the flow harmonics are evaluated. Figure 21 shows $\mathrm{SC}(v_{1}^{2},v_{3}^{2})$ and $\mathrm{SC}(v_{2}^{2},v_{3}^{2})$ for C+Pb collisions, where the cumulants are constructed from ensemble-averaged moments over all particles and events at fixed $N_{\mathrm{part}}$. Both configurations yield a positive $\mathrm{SC}(v_{1}^{2},v_{3}^{2})$, with no statistically significant distinction between them. The $\mathrm{SC}(v_{2}^{2},v_{3}^{2})$ values are consistent with zero for both the $\alpha$-clustering and Woods–Saxon cases.
Figure 21: Symmetric cumulants of $v_{n}$ as a function of $N_{\mathrm{part}}$ for Woods–Saxon and triangular $\alpha$-clustered initial configurations for C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV.
IV Conclusion
In this work, we have investigated the sensitivity of low-energy nuclear collisions to intrinsic nuclear structure by studying C+C and C+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=2.36$ GeV within the JAM transport model. By comparing Woods–Saxon and triangular $\alpha$-clustered configurations for the $^{12}\mathrm{C}$ nucleus, we systematically explored how initial-state geometric features propagate to final-state observables in a regime dominated by baryon stopping, mean-field interactions, and hadronic rescattering.
We find that $\alpha$ clustering leads to a more compact participant configuration than the Woods–Saxon case, while global transverse size and eccentricity fluctuations exhibit only a weak sensitivity to clustering. The root-mean-square radii of the two configurations differ by less than a few percent, indicating that the observed effects are not driven by trivial changes in nuclear size, but rather by differences in the microscopic spatial organization of nucleons. At this beam energy, radial observables remain directly sensitive to such geometric compactness. In particular, the proton mean transverse momentum is enhanced for $\alpha$-clustered configurations, whereas pion observables show little response. The ensemble-averaged $\langle p_{T}\rangle$ fluctuations of protons tend to be smaller in the $\alpha$-clustered configuration than in the Woods–Saxon case, consistent with reduced transverse-radius fluctuations of participant nucleons in the clustered configuration.
The collective anisotropic response of the system is examined through the flow coefficients. We find that $\alpha$ clustering is associated with an enhancement of the root-mean-square flow magnitudes, $v_{n}\{2\}=\sqrt{\langle v_{n}^{2}\rangle}$, at large $N_{\mathrm{part}}$, consistent with modifications of the average initial geometry and its collective response. In contrast, the ensemble-averaged fluctuation strength of individual flow harmonics remains small and is not clearly distinguishable within the present statistical precision. While symmetric cumulants constructed from the initial-state eccentricities exhibit a noticeable sensitivity to clustering, the corresponding ensemble-averaged correlations among final-state flow harmonics do not show a comparably clear separation between clustered and unclustered configurations, suggesting that information on initial-state fluctuations is only partially retained during the dynamical evolution.
Overall, our results demonstrate that low-energy nuclear collisions retain measurable sensitivity to $\alpha$ clustering through a combination of radial observables and correlation-based flow measurements. We note, however, that the $\alpha$-clustered configuration employed in this work represents a simplified geometric implementation, and that more realistic descriptions incorporating state-of-the-art nuclear structure calculations from low-energy nuclear physics will be essential for quantitative constraints on clustering effects. These findings highlight the complementary roles of different classes of observables in probing intrinsic nuclear structure and provide strong motivation for future experimental measurements at CSR and HIAF.
Acknowledgments
SS thanks colleagues at the Institute of Modern Physics for valuable discussions. SS also thanks Dr. Amaresh Jaiswal and Dr. Sandeep Chatterjee for useful discussions during the WHEPP 2025 conference. SS acknowledges support from the Chinese Academy of Sciences.